# Grand-Canonical Symmetric Orbifold Theories

### v1

@@ III.4 Connected Correlation Functions
$\langle\alpha_{1}\cdot\prod_{i}\alpha_{r_{i}}\rangle_{d}^{conn}=N_{c}\cdot\langle\prod_{i}\alpha_{r_{i}}\rangle_{d}^{conn}\,$, (45)
@@ IV Outlook
In this letter, we have shown that grand-canonical ensembles of symmetric orbifold conformal field theories can be formulated in such a way that they satisfy the same axioms as a logarithmic conformal field theory on the sphere. The appropriate formulation is best suited to describe a second-quantized Hilbert space for the string theory. In this context, we have provided a conformal field theory interpretation of the central charge operator of an $AdS_{3}$ string theory at the tensionless point. It would be interesting to understand the field-theoretic origin of the central charge operator beyond the unit-flux case. String theory on an $AdS_{3}\times X$ background with pure NS-NS flux is believed to be dual to a deformed symmetric orbifold in the grand-canonical ensemble [25, 26]. We therefore expect the ideas developed here to be of relevance in the more general case.
Another avenue for future work would be to generalize our construction to symmetric orbifolds with boundaries. This may shed light on some grand-canonical aspects of holographic interfaces in $AdS_{3}/CFT_{2}$ [27, 28, 29, 30]. A first step in this direction was taken in [31], which explores grand-canonical Hurwitz theories with boundaries.
@@ Acknowledgements.
It is a pleasure to thank Jan Troost for many stimulating discussions and for comments on an early version of the manuscript. I am also grateful to António Antunes and Eric Perlmutter for insightful questions.

### v2

@@ III.4 Connected Correlation Functions
$\langle\alpha_{1}\cdot\prod_{i}\alpha_{r_{i}}\rangle_{d}^{\rm conn}=N_{c}\cdot\langle\prod_{i}\alpha_{r_{i}}\rangle_{d}^{\rm conn}\,$, (44)
@@
Finally, we note that on the sphere, disconnected single-sheeted contributions can be factored out of any worldsheet and cancelled against the full partition function. This is akin to the argument used to discard bubble diagrams in quantum field theory. It implies that connected grand-canonical correlators on the sphere reduce to
$\langle\prod_{i}\alpha_{k_{i}}\rangle^{\rm conn}_{p}=\sum_{N_{c}}p^{N_{c}}\cdot\langle\prod_{i}\alpha_{k_{i}}\rangle^{\rm un,\,conn}_{N_{c}}\,$, (45)
where the sum runs over values $N_{c}$ of the degree for which connected covers without extra disconnected spheres exist. The set of allowed values is finite, as it is constrained by the Riemann-Hurwitz formula. This offers yet another perspective on the advantage of going grand-canonical: it eliminates the need to account for disconnected sphere diagrams.
@@ IV Outlook
In this letter, we have shown that grand-canonical ensembles of symmetric orbifold conformal field theories can be formulated as the direct sum of symmetric orbifolds of all degrees. As such, they define unitary and conformally invariant quantum field theories that admit a consistent operator product expansion. Our formulation is best suited to describe a second-quantized Hilbert space for the dual string theory. In this context, we have provided a conformal field theory interpretation of the central charge operator of an $AdS_{3}$ string theory at the tensionless point. It would be interesting to understand the field-theoretic origin of the central charge operator beyond the unit-flux case. String theory on an $AdS_{3}\times X$ background with pure NS-NS flux is believed to be dual to a deformed symmetric orbifold in the grand-canonical ensemble [24, 25]. We therefore expect the ideas developed here to be of relevance in the more general case.
Another avenue for future work would be to generalize our construction to symmetric orbifolds with boundaries. This may shed light on some grand-canonical aspects of holographic interfaces in $AdS_{3}/CFT_{2}$ [26, 27, 28, 29, 30]. A first step in this direction was taken in [31], which explores grand-canonical Hurwitz theories with boundaries.
@@ Acknowledgements.
It is a pleasure to thank Jan Troost for many stimulating discussions and for comments on an early version of the manuscript. I am also grateful to António Antunes, Simon Douaud, Sebastian Harris and Eric Perlmutter for helpful conversations. I thank the anonymous referee for comments that led to improvements and corrections, and DESY for its hospitality during the period in which this work was revised.